OpenStudy (anonymous):

Find the value of: \[\frac{1}{1 + x^{(b-a)} +x^{(c-a)}} + \frac{1}{1+x^{(a-b)} +x^{(c-b)}} + \frac{1}{1+x^{(b-c)}+x^{(a-c)}}\]

OpenStudy (helder_edwin):

why don't u tray \[ \Large u=x^{(a-b)} \] \[ \Large v=x^{(a-c)} \] \[ \Large w=x^{(b-c)} \]

OpenStudy (anonymous):

So, by this it becomes: \[\frac{1}{1 + \frac{1}{u} +v} + \frac{1}{1+u +\frac{1}{w}} + \frac{1}{1+w+v} \] Then what??

OpenStudy (anonymous):